Now consider an economy consists of four workers: Harry, Larry, Moe, and Curly, where Harry and Larry are brothers. Each works 10 hours a day and can produce two services: mowing lawns and washing cars. In an hour, both Harry and Larry can either mow one lawn or wash one car; Moe can either mow one lawn or wash two cars; and Curly can either mow two lawns or wash one car.

g. If there are 6 heterogeneous workers, how many kink points would be there on the PPF?
h. If there are 6 homogeneous works, how many kink points would be there on the PPF?
i. If there are 6 workers, how many kink points would be there on the PPF?
j. If there are 6 workers, these workers could be divided into 2 categories (high or low) according to
their skills, how many possible labor force compositions?
k. If there are N heterogeneous workers, how many kink points would be there on the PPF?
l. If N infinite, what would happen to the PPF?
1
Expert's answer
2019-10-02T09:27:23-0400

g. If there are 6 heterogeneous workers, then there would be up to 6×6 = 36 kink points there on the PPF.

h. If there are 6 homogeneous works, there would no kink points there on the PPF.

i. If there are 6 workers of 2 groups, then there would be up to 6×6 = 36 kink points there on the PPF.

j. If there are 6 workers, these workers could be divided into 2 categories (high or low) according to

their skills, then there would be up to 6×2 = 12 kink points there on the PPF.

k. If there are N heterogeneous workers, then there would be up to N^2 kink points there on the PPF.

l. If N infinite, then the PPF would become a line.